Use the aggregate expenditures model and assume the marginal propensity to consume is 0.90. An increase in government spending of $1 billion would result in an increase in GDP of?
Answer #1

Answer

Multiplier = 1/(1 - MPC)

where MPC = Marginal propensity to consume = 0.9

=> Multiplier = 1/(1 - MPC) = 1/(1 - 0.9) = 10

This means that increase in autonomous expenditure by $1 will result in increase in GDP by $10. Here government spending increases by $1 billion will result in increase in Autonomous expenditure by 1 billion.

Hence This increase in autonomous expenditure will increase GDP by 10*1 billion = 10 billion.

Hence, This increase in government spending will leads to increase in GDP by $10 billion.